In testing the hypotheses H​0: p​ = 0.65 vs Ha​: p​ =/= 0.65, the test statistic is calculated to be z​ = 2.07. Which of the following is the correct​ p-value?
Answer #1

this is two tailed test

z= 2.07

P(z > 2.07) = 1 - P(z <  2.07) = 1-0.981=0.019*2=0.038

P-value = 0.038( using z table )

-------------------------------------

DEAR STUDENT,

IF YOU HAVE ANY QUERY ASK ME IN THE COMMENT BOX,I AM HERE TO HELPS YOU.PLEASE GIVE ME POSITIVE RATINGS

*****************THANK YOU***************